OpenStudy (anonymous):

identify the slope and y-intercept in 3y+6x=-9 and how to graph it

OpenStudy (cwrw238):

convert to the slope - interccept for y = mx + c

OpenStudy (jdoe0001):

aka, solve for "y"

OpenStudy (anonymous):

Do you need help with like @jdoe0001 said solving for y or graphing?

OpenStudy (anonymous):

@radum ??

OpenStudy (anonymous):

3y+6x=-9 so, subtract 3y from both sides 6x=-3y-9 then, add 9 to both sides 6x+9= 3y what do u think happens next ?

OpenStudy (anonymous):

hint "isolate y", so to undo 3y divide. can u show me the next step??

OpenStudy (anonymous):

I guess u understand it?